# ProFit: Leveraging High-Value Signals in SFT via Probability-Guided Token Selection

### 25 Mar 2026 v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable general capabilities (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025a). To adapt them to specific downstream tasks, Supervised Fine-Tuning (SFT) has become the prevailing paradigm (Chung et al., 2024). Traditional SFT is based on an autoregressive objective, forcing the model to strictly align with a single reference answer at the token level. However, this rigid objective neglects the one-to-many nature of language (Li et al., 2016; Yang et al., 2025b), where diverse expressions can convey the same intent. Therefore, this strategy of forcibly fitting a single reference is often suboptimal and can easily lead to the model simply memorizing specific samples. (Gudibande et al., 2023; Chu et al., 2025).
@@ 2.1 Data-Efficient Instruction Tuning
Recent SFT adheres to the less is more principle (Zhou et al., 2023a; Zhang et al., 2025; Li et al., 2025c), evolving from complexity-based selection (Cao et al., 2023; Li et al., 2024a) to 2025’s importance-aware metrics like MIWV (Jiang et al., 2025) and ICL-based filtering (Wang et al., 2025a; Jiang et al., 2025). However, these coarse-grained sample-level methods treat pairs as atomic units, overlooking intra-sample low-information segments or stylistic noise (Pang et al., 2025; Qin et al., 2025), thus limiting the model’s focus on dense logical signals.
@@ 2.2 Token-Level Training Objectives
To address granularity limitations, research has pivoted to token-level optimization. While classical methods like Focal Loss (Lin et al., 2017) and Unlikelihood Training (Welleck et al., 2019) targeted hard tokens, recent LLM approaches like Rho-1 (Lin et al., 2024) and TIS-DPO (Liu et al., 2024) rely on costly external reference models. More efficient intrinsic methods like DFT (Wu et al., 2025b) employ probability-driven soft reweighting, and CFT (Ruan et al., 2025) further validates the need for supervising critical regions. Recently, Li et al. (2025a) theoretically validated via a model-capability continuum that suppressing low-probability tokens benefits strong-prior domains like math, supporting our masking strategy. While sharing the probability-driven motivation with DFT, ProFit diverges by adopting a strict hard masking strategy to efficiently filter out non-core expressions in a single step.
@@ 5.1 Experimental Setup
  • Entropy-based tuning (Wang et al., 2025b): A selective strategy that updates parameters only on tokens with high entropy.
@@ 5.2 Main Results
In summary, ProFit shows strong scalability and universality, consistently boosting performance across model sizes (0.6B–14B) and architectures without the high data costs of multi-reference fine-tuning.
@@ 7 Conclusion
In this work, we propose ProFit, a novel method to fine-tune LLMs by leveraging token prediction probabilities as a proxy for semantic importance. Inspired by our hypothesis testing results revealing that high-probability tokens carry core semantics while low-probability ones correspond to non-core expressions, we propose ProFit to selectively mask the latter, aiming to alleviate the overfitting to surface-level phrasing and capture the underlying logic. Extensive experiments across multiple LLM series, including Qwen, Llama, and OLMo2, demonstrate that ProFit consistently outperforms conventional full-parameter SFT and other data selection methods on diverse benchmarks covering general reasoning, mathematics, and instruction following. We further conduct comprehensive ablation studies on training epochs, probability thresholds, and LoRA ranks. These analyses confirm that low-probability tokens serve as a primary source of optimization interference, validating that our approach offers a robust and effective solution for improving model generalization.

### 06 May 2026 v3

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable general capabilities (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025a). To adapt them to specific downstream tasks, Supervised Fine-Tuning (SFT) has become the prevailing paradigm (Chung et al., 2024). Traditional SFT is based on an autoregressive objective, forcing the model to strictly align with a single reference answer at the token level. However, this rigid objective neglects the one-to-many nature of language (Li et al., 2016; Yang et al., 2025b), where diverse expressions can convey the same intent. Therefore, this strategy of forcibly fitting a single reference is often suboptimal and can easily lead to the model simply memorizing specific samples. (Gudibande et al., 2023; Chu et al., 2025; Wang et al., 2025a, 2026).
@@ 2.1 Data-Efficient Instruction Tuning
Recent SFT adheres to the less is more principle (Zhou et al., 2023a; Zhang et al., 2025; Li et al., 2025c; Ji et al., 2026), evolving from complexity-based selection (Cao et al., 2023; Li et al., 2024a) to 2025’s importance-aware metrics like MIWV (Jiang et al., 2025) and ICL-based filtering (Wang et al., 2025b; Jiang et al., 2025). However, these coarse-grained sample-level methods treat pairs as atomic units, overlooking intra-sample low-information segments or stylistic noise (Pang et al., 2025; Qin et al., 2025), thus limiting the model’s focus on dense logical signals.
@@ 2.2 Token-Level Training Objectives
To address granularity limitations, research has pivoted to token-level optimization (Kong et al., 2025). While classical methods like Focal Loss (Lin et al., 2017) and Unlikelihood Training (Welleck et al., 2019) targeted hard tokens, recent LLM approaches like Rho-1 (Lin et al., 2024) and TIS-DPO (Liu et al., 2024) rely on costly external reference models. More efficient intrinsic methods like DFT (Wu et al., 2025b) employ probability-driven soft reweighting, and CFT (Ruan et al., 2025) further validates the need for supervising critical regions. Recently, Li et al. (2025a) theoretically validated via a model-capability continuum that suppressing low-probability tokens benefits strong-prior domains like math, supporting our masking strategy. While sharing the probability-driven motivation with DFT, ProFit diverges by adopting a strict hard masking strategy to efficiently filter out non-core expressions in a single step.
@@ 5.1 Experimental Setup
  • Entropy-based tuning (Wang et al., 2025c): A selective strategy that updates parameters only on tokens with high entropy.
@@ 5.2 Main Results
In summary, ProFit shows strong scalability and universality, consistently boosting performance across model sizes and architectures without the high data costs of multi-reference fine-tuning.
@@ 6.4 Superior RL Initialization
6.5 Analysis of Fine-Tuning Instruct Models
Further fine-tuning instruction-aligned models typically incurs an alignment tax, degrading general capabilities Wu et al. (2025a, b); Diao et al. (2026). We evaluate ProFit on the Qwen2.5-Instruct series to test its robustness against this issue. As shown in Table 2, while standard SFT and DFT consistently suffer performance drops (e.g., -1.83 average degradation for SFT on 7B), ProFit effectively mitigates catastrophic forgetting. It not only prevents typical declines but actually improves the average scores for the 0.5B (+1.24) and 7B (+0.17) models, with notable gains in reasoning tasks like GSM8K and AIME’24, proving its superiority in refining aligned models without sacrificing foundational instruction-following skills.
Model Method GPQA-Diamond GSM8K MATH-500 IFEval AIME’24 Avg. $\Delta$
Qwen2.5-0.5B-Instruct Vanilla 17.42 37.38 27.98 25.28 0.21 21.65 -
SFT 17.55 37.96 24.47 23.66 0.21 20.77 -0.88
DFT 10.67 40.48 27.65 25.25 0.31 20.87 -0.78
[…]
Qwen2.5-3B-Instruct Vanilla 32.01 74.73 65.78 58.99 4.38 47.18 -
SFT 30.30 75.29 63.52 57.44 5.31 46.37 -0.81
DFT 26.58 77.27 65.77 50.72 7.60 45.59 -1.59
[…]
Qwen2.5-7B-Instruct Vanilla 35.48 81.00 75.28 69.29 12.50 54.71 -
SFT 33.27 83.82 74.65 61.95 10.73 52.88 -1.83
DFT 32.77 85.60 74.65 60.28 13.02 53.26 -1.45
[…]
Table 2: Performance evaluation on previously instruction-tuned models (Qwen2.5-Instruct series). We report the accuracy of the original Instruct model (Vanilla), standard SFT, DFT, and ProFit.
Model Method HellaSwag ARC-c HumanEval MMLU-Pro History MMLU-Pro Chemistry Avg. $\Delta$
Qwen3-0.6B-Base Vanilla 24.54 1.69 1.45 1.31 3.18 6.43 -
SFT 33.37 45.76 32.09 15.22 15.72 28.43 +22.00
Entropy 33.16 50.51 31.78 11.02 15.90 28.47 +22.04
DFT 32.72 53.22 37.80 10.76 14.93 29.89 +23.46
[…]
Qwen3-4B-Base Vanilla 65.55 34.92 75.38 27.03 16.96 43.97 -
SFT 74.83 92.20 82.85 41.99 62.46 70.87 +26.90
Entropy 74.77 90.85 81.78 41.73 61.31 70.09 +26.12
DFT 62.12 90.85 81.40 42.26 59.28 67.18 +23.21
[…]
Qwen3-14B-Base Vanilla 84.02 94.92 83.69 54.59 57.95 75.03 -
SFT 82.18 94.58 86.97 53.81 70.05 77.52 +2.49
Entropy 81.43 93.90 87.04 55.38 70.41 77.63 +2.60
DFT 86.96 94.58 87.65 56.43 69.70 79.06 +4.03
[…]
Table 3: Performance comparison on General Tasks across different Qwen3 model scales (0.6B, 4B, 14B). We report the accuracy of the Vanilla baseline, standard SFT, and varying fine-tuning strategies (Entropy, DFT, ProFit).
6.6 Analysis of Performance on General Tasks
To evaluate whether ProFit preserves comprehensive abilities while optimizing for specific domains, we tested it on general benchmarks across Qwen3 scales. As shown in Table 3, ProFit demonstrates strong generalization without severe catastrophic forgetting. It achieves the highest average improvements on the 0.6B (+26.56) and 14B scales over the Vanilla model. Even on the 4B scale, where standard SFT slightly edges out in the overall average, ProFit remains highly competitive and excels in specific reasoning-heavy subsets , proving it effectively balances broad knowledge retention with its primary fine-tuning objectives.
@@ 7 Conclusion
In this work, we propose ProFit, a novel method to fine-tune LLMs by leveraging token prediction probabilities as a proxy for semantic importance. Inspired by our hypothesis testing results revealing that high-probability tokens carry core semantics while low-probability ones correspond to non-core expressions, we propose ProFit to selectively mask the latter, aiming to alleviate the overfitting to surface-level phrasing and capture the underlying logic. Extensive experiments across multiple LLM series, including Qwen, Llama, and OLMo2, demonstrate that ProFit consistently outperforms conventional full-parameter SFT and other data selection methods on diverse benchmarks covering general reasoning, mathematics, and instruction following. We further conduct comprehensive ablation studies on training epochs, probability thresholds, and LoRA ranks. ProFit also serves as a superior initialization for reinforcement learning, promoting stable convergence and deeper reasoning exploration. These analyses confirm that low-probability tokens serve as a primary source of optimization interference, validating that our approach offers a robust and effective solution for improving model generalization.
Acknowledgements
This work was partly supported by the National Natural Science Foundation of China (Grant No. 62576191), the Shenzhen Science and Technology Program (ZDCY20250901103533010) and Tsinghua SIGS KA Cooperation Fund.
@@ Appendix F Benchmark Details
  • HellaSwag (Zellers et al., 2019): A benchmark for evaluating commonsense natural language inference, requiring models to choose the most logical and natural continuation of a given text snippet.
  • ARC-c (Clark et al., 2018): The Challenge set of the AI2 Reasoning Challenge, consisting of difficult grade-school science questions that require advanced logic and reasoning beyond simple information retrieval.
  • HumanEval (Chen et al., 2021): A benchmark designed to evaluate code generation capabilities, featuring programming problems (primarily in Python) that test algorithmic logic and basic coding skills.
  • MMLU-Pro (Wang et al., 2024): An enhanced, more challenging version of the Massive Multitask Language Understanding (MMLU) benchmark. In our evaluation, we specifically report performance on the History and Chemistry subsets to assess advanced, domain-specific knowledge and reasoning.